# Dualtronics: leveraging both faces of polar semiconductors

### v1

@@ Acknowledgements
Len van Deurzen and Eungkyun Kim contributed equally to this work.
The work at the Institute of High Pressure Physics PAS was supported by the National Science Centre (NCN) of Poland, Grant No. 2021/43/D/ST3/03266 and European Union under HORIZON EUROPE Digital, Industry and Space (VISSION ID:101070622),
This work was performed in part at the Cornell NanoScale Facility, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation (Grant NNCI-2025233). This work further made use of the electron microscopy facility of the Cornell Center for Materials Research (CCMR) with support from the National Science Foundation Materials Research Science and Engineering Centers (MRSEC) program (DMR1719875). The Thermo Fisher Spectra 300 X-CFEG was acquired with support from PARADIM, an NSF MIP (DMR-2039380) and Cornell University.

### v2

@@ Acknowledgements
The work at the Institute of High Pressure Physics PAS was supported by the National Science Centre (NCN) of Poland with Grant No. 2021/43/D/ST3/03266 and the European Union under HORIZON EUROPE Digital, Industry and Space (VISSION ID:101070622).
This work was performed in part at the Cornell NanoScale Facility, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation (Grant NNCI-2025233). This work further made use of the electron microscopy facility of the Cornell Center for Materials Research (CCMR) with support from the National Science Foundation Materials Research Science and Engineering Centers (MRSEC) program (DMR1719875). The Thermo Fisher Spectra 300 X-CFEG was acquired with support from PARADIM, an NSF MIP (DMR-2039380) and Cornell University. Further funding was provided by the Army Research Office (ARO) with Grant No. W911NF-22-2-0177 and the National Science Foundation with Grant No. DGE-2139899. The authors thank Seohong Park for illustrating Figure 3(e).
Author contributions
L.v.D. and E.K. contributed equally. L.v.D. and E.K. were responsible for the process development, device fabrication, device characterization, and monolithic switching measurements. H.T. designed the epistructures and developed the growth process and sequence. M.S. and M.C. were in charge of the HEMT and LED MBE growth. A.F-Z. prepared the substrate and performed atomic force microscopy. Z.Z. carried out the MBE regrowth of the HEMT source and drain contacts, and performed atomic force microscopy. N.P. and D.M. provided STEM images. D.J., L.v.D., E.K., Z.Z., H.G.X. and H.T. developed the scientific and technological objectives. L.v.D. and E.K. prepared the manuscript. D.J and H.T. revised the manuscript. D.J., H.G.X., and H.T. oversaw the project.
Competing interests
The authors filed a US provisional patent (63/548,460) based on the design and technology laid out in this article.
Additional information
For correspondence and for questions regarding the materials, L.v.D., E.K., D.J. and H.T. may be contacted.